Given that oil, with a density p = 700 kg/m3 and Kinematic viscosity v = 0.00001m2/s flows at 0.3m3/s through 500m of 250mm diameter cast iron pipe and the roughness of cast iron pipe e = 0.26mm


Take note that the pipe runs at an angle of 10ndegrees


  • Determine the Reynolds number
  • The friction factor f
  • The head loss due to friction
1
Expert's answer
2021-04-12T01:55:22-0400

Velocity of flow, V = Q/A = Q/("\pi"/4 x D2) = 0.3 m3/s / (3.14/4 x (0.25 m)2) = 6.11 m/s

Reynolds number, Re = VD/"\upsilon" = (6.11 m/s x 0.25 m) / 0.00001 m2/s = 152 750

The Reynolds number is greater than 2000. So the flow is turbulent. 

Friction factor for turbulent flow, f = [1.14 + 2log(D/e)]-2 = [1.14 + 2log(0.25/0.00026)]-2 = 0.0198

Head loss, HL = f(L/D)(V2/2g) = 0.0198(500 m/0.25 m)((6.11 m/s)2/(2 x 9.81 m/s2) = 75.35 m